OpenStudy (anonymous):

If a 100 milligram tablet of an asthma drug is taken orally and if none of the drug is in the body when the tablet if first taken, the total amount "A" in the bloodstream after "t" minutes is predicted to be. A=100[1-(0.9)^t]; for 0

OpenStudy (anonymous):

6.579mins

OpenStudy (anonymous):

Thanks! can you explain how to solve?? No clue - book is confusing me.. . .

OpenStudy (anonymous):

A=100(1-(.9)^t) A=50 50=100(1-(.9)^t) 1/2=1-(.9)^t -1/2=-(.9)^t 1/2=(.9)^t from here u have to use logs. I'm assuming ur learning about them now?

OpenStudy (anonymous):

u assumed correctly!! thanks for making this easily understood. . .

OpenStudy (anonymous):

ok so u take the log of both sides im gonna use ln here (the natural log) just because i like it ln(1/2)=ln((.9)^t) now the rules of logs is that if u have an exponent inside the log u can take it out and multiply times the whole log so ln((.9)^t) becomes t*ln(.9) so ln(1/2)=t*ln(.9) so t=ln(1/2)/ln(.9)

OpenStudy (anonymous):

Your a lifesaver. thanks